OpenStudy (anonymous):

Hey how do you do the definition of sqrt(3x+1)??

OpenStudy (anonymous):

\[\sqrt{3x}+\sqrt{1}\]

OpenStudy (anonymous):

is this what you are looking for?

OpenStudy (atlas):

@rrp240 you are wrong @MathGeekinProgress : your question isn't clear to me

OpenStudy (anonymous):

No I'm supposed to put this into [f(x+h) - f(x)] all over h aka definition form

OpenStudy (psymon):

Definition of a derivative we mean then

OpenStudy (atlas):

definition -> derivation

OpenStudy (atlas):

:P

OpenStudy (psymon):

Well, we literally would replace every x in our function with x+h and then just write it out as the formula describes: \[\frac{ \sqrt{3(x+h)+1} -\sqrt{3x+1}}{ h }\] From here, you would be forced to multiply top and bottom by the conjugate, which is eseentially taking the whole numerator and changing the middle sign into a plus. So it would look like this: \[\frac{ \sqrt{3(x+h)+1} - \sqrt{3x+1} }{ h }*\frac{ \sqrt{3(x+h)+1} + \sqrt{3x+1} }{ \sqrt{3(x+h)+1} + \sqrt{3x+1} }\]

OpenStudy (anonymous):

Ok

OpenStudy (psymon):

If you can multiply out and simplify the top properly, it shouldnt be too bad. It is messy, so maybe as a tip to try and multiply it out safely (only a suggestion), know that: \[(a+b)(a-b) = a^{2} - b^{2}\]If you undersrand that, it makes it easy to foil out the top without much trouble.

OpenStudy (anonymous):

So the sqrt symbols on the top will go away since we are basically squaring it. So we would come to this: \[(3h + 2) / (h(\sqrt{3x+3h+1} - \sqrt{3x+1})\] right?